### v1

@@ Abstract
Recent advances in LiDAR 3D detection have demonstrated the effectiveness of Transformer-based frameworks in capturing the global dependencies from point cloud spaces, which serialize the 3D voxels into the flattened 1D sequence for iterative self-attention. However, the spatial structure of 3D voxels will be inevitably destroyed during the serialization process. Besides, due to the considerable number of 3D voxels and quadratic complexity of Transformers, multiple sequences are grouped before feeding to Transformers, leading to a limited receptive field. Inspired by the impressive performance of State Space Models (SSM) achieved in the field of 2D vision tasks, in this paper, we propose a novel Unified Mamba (UniMamba), which seamlessly integrates the merits of 3D convolution and SSM in a concise multi-head manner, aiming to perform “local and global” spatial context aggregation efficiently and simultaneously. Specifically, a UniMamba block is designed which mainly consists of spatial locality modeling, complementary Z-order serialization and local-global sequential aggregator. The spatial locality modeling module integrates 3D submanifold convolution to capture the dynamic spatial position embedding before serialization. Then the efficient Z-order curve is adopted for serialization both horizontally and vertically. Furthermore, the local-global sequential aggregator adopts the channel grouping strategy to efficiently encode both “local and global” spatial inter-dependencies using multi-head SSM. Additionally, an encoder-decoder architecture with stacked UniMamba blocks is formed to facilitate multi-scale spatial learning hierarchically. Extensive experiments are conducted on three popular datasets: nuScenes, Waymo and Argoverse 2. Particularly, our UniMamba achieves 70.2 mAP on the nuScenes dataset.
@@ 1 Introduction
Voxel-based methods can be further categorized into sparse convolution neural networks (SpCNN)-based and Transformer-based. SpCNN-based approaches [46, 44] are constrained by the inherently small receptive field of convolutions, which limits their ability to extract rich contextual information. Transformer-based methods [24, 21, 37] suffer from quadratic complexity, which tend to serialize 3D voxels into multiple 1D sequences and adopt self-attention mechanism respectively, leading to limited receptive field by group size and feature size (Fig. 1 (a)). However, both local structure details and global context are crucial for 3D object detection from sparse and occluded point clouds. Under this circumstance, a robust LiDAR 3D backbone with flexible receptive fields and great spatial modeling capacity as well as computational efficiency is worth exploring.
Inspired by the remarkable success of the Mamba [11] architecture in the areas of natural language processing and 2D vision, we investigate its applicability to LiDAR 3D detection. Its linear complexity enables the whole sequence processing without voxel grouping (Fig. 1 (b)), thus facilitating the global context modeling. However, simple adaptation of Mamba to LiDAR 3D backbone may have some drawbacks: (1) loss of spatial voxel locality during serialization; (2) local redundancy brought by coherent global modeling, causing inferior efficiency; (3) lack of spatial diversity to handle complex local and global dependencies.
@@ 2.3 Vision Task with Mamba
Mamba [11] is a popular sequential modeling architecture with linear complexity. Due to its great computational efficiency, it has been regarded as a strong competitor to Transformer. Some research has begun to explore its potential in vision tasks [20, 16]. As a sequence model, most existing studies have focused on exploring different scanning methods. Vim [51] introduced a bidirectional SSM and position embeddings to learn image features. GrootVL [40] adopted a tree topology to construct scanning sequences, enhancing long-range interactions. In this paper, we investigate the feasibility of Mamba-style architecture for LiDAR 3D detection, and propose UniMamba, which can effectively capture local and global voxel interactions simultaneously.
@@ 3.3.1 Spatial Locality Modeling
Like other Mamba-based architectures [20, 16, 12], compressing 2D/3D space into a 1D sequence inevitably results in the loss of local spatial position information. To address this issue, current methods often design various complex serialization techniques to maintain locality. However, these approaches tend to incur high computational costs and yield only limited effectiveness. We tackle this problem effectively by introducing Spatial Locality Modeling module (SLM) with a simple submanifold convolution:
@@ 3.3.2 Complementary Z-order Serialization
The purpose of serializing non-empty voxels is to find a path that traverses all input voxels. To preserve spatial topology, the serialization of 3D voxels can be viewed as a space-filling curve process. Existing space-filling curves mainly include Hilbert [14] and Z-order (Morton) [25] curves. While the Hilbert curve offers a better locality, constructing its index is computationally expensive. Thanks to the locality enhancement design (see Sec. 3.3.1) of UniMamba, we do not heavily rely on preserving spatial topology information during serialization. Therefore, we opt for the computationally efficient Z-order curve. Specifically, given the coordinates of the input voxel set $\mathcal{C}=\{(x_{i},y_{i},z_{i}),i\in(1,N)\}$, where N denotes the number of voxels, the corresponding Z-order code $\mathcal{Z}^{P}=\{z_{i}^{P},i\in(1,N)\}$ is computed by the bit-interleaving function [25]. Finally, sorting the Z-order code $\mathcal{Z}^{P}$ in ascending order to gain the Z-order index $\mathcal{Z}^{index}=\{z_{i}^{index},i\in(1,N)\}$ for each voxel.
Traditional Z-order curve using the X-axis as the primary order to calculate the index. However, we observe that indexing solely in the X-direction results in the preservation of spatial proximity only along the X-axis. Therefore, we introduce a variant that uses the Y-axis as the primary order for indexing. This further ensures that voxels that are close in the Y-direction maintain proximity, effectively enhancing perceptual capabilities (see Tab. 4).
@@ 3.3.3 Local-Global Sequential Aggregator
Global Sequential Encoder (GSE) treats all non-empty voxels in the scene as a 1D sequence to directly establish a global receptive field. Unlike traditional window-based serialization, it does not create explicit local windows. The basic GSE layer includes two cascaded Mamba layers to encode the serialized voxels using Complementary Z-order. Inspired by Vim [51], we use a bidirectional SSM as our base Mamba Layer. Specifically, given the input voxel sets $\mathcal{F}\in\mathbb{R}^{N\times C}$ and their corresponding voxel coordinates sets $\mathcal{C}\in\mathbb{R}^{N\times 3}$, We first serialize $\mathcal{F}$ using X-axis Z-order to obtain the corresponding voxel sequence $\mathcal{F}^{X}$ and coordinates $\mathcal{C}^{X}$. Subsequently, the reordered sequence $\mathcal{F}^{X}$ is input into the first Mamba layer to obtain the output $\mathcal{F}^{X}_{out}$. After that, the sequence $\mathcal{F}_{out}^{X}$ is reordered using the Y-axis Z-order curve to obtain the voxel sequence $\mathcal{F}^{Y}$ and coordinates $\mathcal{C}^{Y}$ and input into the second Mamba layer to produce the output $\mathcal{F}^{G}$ of the GSE. The specific calculations can be summarized as follows:
@@ 3.4 The UniMamba 3D Backbone
With the proposed UniMamba Block, we build UniMamba 3D Backbone with flexible receptive fields and great spatial modeling capacity. The architecture of UniMamba is shown in Fig. 2. Following [45], we use an encoder-decoder architecture to stack UniMamba Blocks for further hierarchical feature extraction. In each stage, the number of UniMamba Blocks at each scale is set to 1. As a general voxel-based 3D backbone, we replace the backbone in DSVT [37] to construct our detector. For Argoverse 2 [38], we build on SAFDNet [46]. The configuration of the detection head and loss functions remains consistent with the baseline.
@@ 4.1 Datasets and Metrics
nuScenes [2] is a sophisticated outdoor dataset comprising 1,000 scenarios, with 700 allocated for training, 150 for validation, and 150 for testing. It encompasses a diverse array of object annotations across 10 categories, representing the majority of traffic participants. The size of objects varies significantly among these categories, which poses considerable challenges for detection. We evaluate the effectiveness of our approach using the provided assessment tools, and reporting metrics including NDS and mAP. The NDS takes into account average detection accuracy as well as other error factors(e.g., direction, speed).
@@ 4.2 Implementation Details
We implement our UniMamba based on OpenPCDet [35]. Following previous methods [46], we use voxel sizes of $(0.3m,0.3m,0.25m)$, $(0.32m,0.32m,0.1875m)$, and $(0.4m,0.4m,0.25m)$ for voxelization of the nuScenes, Waymo, and Argoverse 2 datasets, respectively, where the number of voxel feature channels is set to 128. The downsampling stride for encoder-decoder is set to $\{1,2,2\}$, with each scale containing 1 layer of UniMamba Block. For each UniMamba block, the number of channel groups $M$ is set to 4, with $J=2$ groups applying GSE and the remaining two applying LSE. We trained for 48 epochs on nuScenes without CBGS [50], 24 epochs on Waymo, and 12 epochs on the Argoverse 2 dataset. All experiments are conducted using 8 Tesla A800 GPUs, utilizing the AdamW [23] optimizer. For more network details, please refer to the appendix.
@@ 4.4 Ablation Study
Effect of Spatial Locality Modeling. To validate the effectiveness of the proposed SLM in mitigating the loss of spatial locality, we compare it with several popular space-filling curves. The ”Random” refers to the configuration where no spatial curve is applied. To ensure the same parameter count as Complementary Z-order Serialization, all experiments use a GSE-only structure with two bidirectional SSM layers. As shown in rows 2-4 of Tab. 4, when neither SLM nor any spatial curve is used, performance drops significantly (- 1.4 mAP), demonstrating the importance of spatial locality. Compared with the standard Z-order curve, the Hilbert curve which better preserves locality shows a 0.2 mAP advantage. With SLM, the ”Random” achieves 67.4 mAP, comparable to the performance with space-filling curves. Additionally, the Z-order curve achieves performance parity with the Hilbert curve (67.6 mAP) but requires only 1/17 time cost. These results demonstrate the effectiveness of Spatial Locality Modeling in capturing the dynamic structure embedding.

### v2

@@ Abstract
Recent advances in LiDAR 3D detection have demonstrated the effectiveness of Transformer-based frameworks in capturing the global dependencies from point cloud spaces, which serialize the 3D voxels into the flattened 1D sequence for iterative self-attention. However, the spatial structure of 3D voxels will be inevitably destroyed during the serialization process. Besides, due to the considerable number of 3D voxels and quadratic complexity of Transformers, multiple sequences are grouped before feeding to Transformers, leading to a limited receptive field. Inspired by the impressive performance of State Space Models (SSM), in this paper, we propose a novel Unified Mamba (UniMamba), which seamlessly integrates the merits of 3D convolution and SSM in a concise multi-head manner, aiming to perform “local and global” spatial context aggregation efficiently and simultaneously. Specifically, a UniMamba block is designed which mainly consists of spatial locality modeling, complementary Z-order serialization and local-global sequential aggregator. The spatial locality modeling module integrates 3D submanifold convolution to capture the dynamic spatial position embedding before serialization. Then the efficient Z-order curve is adopted for serialization both horizontally and vertically. Furthermore, the local-global sequential aggregator adopts the channel grouping strategy to efficiently encode both “local and global” spatial inter-dependencies using multi-head SSM. Additionally, an encoder-decoder architecture with stacked UniMamba blocks is formed to facilitate multi-scale spatial learning hierarchically. Extensive experiments are conducted on three popular datasets: nuScenes, Waymo and Argoverse 2. Particularly, our UniMamba achieves 70.2 mAP on the nuScenes dataset.
@@ 1 Introduction
Voxel-based methods can be further categorized into sparse convolution neural networks (SpCNN)-based and Transformer-based. SpCNN-based approaches [49, 47] are constrained by the inherently small receptive field of convolutions, which limits their ability to extract rich contextual information. Transformer-based methods [27, 23, 40] suffer from quadratic complexity, which tend to serialize 3D voxels into multiple 1D sequences and adopt self-attention mechanism respectively, leading to limited receptive field by group size and feature size (Fig. 1 (a)). However, both local structure details and global context are crucial [18, 19, 31] for 3D object detection from sparse and occluded point clouds. Under this circumstance, a robust LiDAR 3D backbone with flexible receptive fields and great spatial modeling capacity as well as computational efficiency is worth exploring.
Inspired by the remarkable success of the Mamba [10, 54, 24, 22] architecture in the areas of vision tasks, we investigate its applicability to LiDAR 3D detection. Its linear complexity enables the whole sequence processing without voxel grouping (Fig. 1 (b)), thus facilitating the global context modeling [50]. However, simple adaptation of Mamba to LiDAR 3D backbone may have some drawbacks: (1) loss of spatial voxel locality during serialization; (2) local redundancy brought by coherent global modeling, causing inferior efficiency; (3) lack of spatial diversity to handle complex local and global dependencies.
@@ 2.3 Vision Task with Mamba
Mamba [10] is a popular sequential modeling architecture with linear complexity. Due to its great computational efficiency, it has been regarded as a strong competitor to Transformer. Some research has begun to explore its potential in vision tasks [22, 15, 20, 11, 21]. As a sequence model, most existing studies have focused on exploring different scanning methods. Vim [54] introduced a bidirectional SSM to learn image features. GrootVL [43] adopted a tree topology to construct scanning sequences, enhancing long-range interactions. The most advanced lidar detectors utilized the Mamba architecture, while LION [24] developed a unified RNN framework using group-based strategy. VoxelMamba [50] introduced a group-free approach to directly model global information. In this paper, we investigate the feasibility of Mamba-style architecture for LiDAR 3D detection, and propose UniMamba, which can effectively capture local and global voxel interactions simultaneously.
@@ 3.3.1 Spatial Locality Modeling
Like other Mamba-based architectures [22, 15, 11, 50], compressing 2D/3D space into a 1D sequence inevitably results in the loss of local spatial position information. To address this issue, most methods often design various complex serialization techniques to maintain locality. However, these approaches tend to incur high computational costs and yield only limited effectiveness. Inspired by previous works [24, 42] that use sparse convolutions to provide positional information, we introduce a simple submanifold convolution as Spatial Locality Modeling module (SLM) within the UniMamba block:
@@ 3.3.2 Complementary Z-order Serialization
The purpose of serializing non-empty voxels is to find a path that traverses all input voxels. To preserve spatial topology, the serialization of 3D voxels can be viewed as a space-filling curve process. Existing space-filling curves mainly include Hilbert [13] and Z-order  [28] curves. While the Hilbert curve offers a better locality, constructing its index is computationally expensive. Thanks to the locality enhancement design (see Sec. 3.3.1), we do not heavily rely on preserving spatial topology information during serialization. Therefore, we opt for the computationally efficient Z-order curve. Specifically, given the coordinates of the input voxel set $\mathcal{C}=\{(x_{i},y_{i},z_{i}),i\in(1,N)\}$, where N denotes the number of voxels, the corresponding Z-order code $\mathcal{Z}^{P}=\{z_{i}^{P},i\in(1,N)\}$ is computed by the bit-interleaving function [28]. Finally, sorting the Z-order code $\mathcal{Z}^{P}$ in ascending order to gain the Z-order index $\mathcal{Z}^{index}=\{z_{i}^{index},i\in(1,N)\}$ for each voxel.
Traditional Z-order curve using the X-axis as the primary order to calculate the index. However, we observe that indexing solely in the X-direction results in the preservation of spatial proximity only along the X-axis. Therefore, we introduce a variant that uses the Y-axis as the primary order for indexing.
@@ 3.3.3 Local-Global Sequential Aggregator
Global Sequential Encoder (GSE) treats all non-empty voxels in the scene as a 1D sequence to directly establish a global receptive field. Unlike traditional window-based serialization, it does not create explicit local windows. The basic GSE layer includes two cascaded Mamba layers to encode the serialized voxels using Complementary Z-order. Specifically, given the input voxel sets $\mathcal{F}\in\mathbb{R}^{N\times C}$ and their corresponding voxel coordinates sets $\mathcal{C}\in\mathbb{R}^{N\times 3}$, We first serialize $\mathcal{F}$ using X-axis Z-order to obtain the corresponding voxel sequence $\mathcal{F}^{X}$ and coordinates $\mathcal{C}^{X}$. Subsequently, the reordered sequence $\mathcal{F}^{X}$ is input into the first Mamba layer to obtain the output $\mathcal{F}^{X}_{out}$. After that, the sequence $\mathcal{F}_{out}^{X}$ is reordered using the Y-axis Z-order curve to obtain the voxel sequence $\mathcal{F}^{Y}$ and coordinates $\mathcal{C}^{Y}$ and input into the second Mamba layer to produce the output $\mathcal{F}^{G}$ of the GSE. The specific calculations can be summarized as follows:
@@ 3.4 The UniMamba 3D Backbone
With the proposed UniMamba Block, we build UniMamba 3D Backbone with flexible receptive fields and great spatial modeling capacity. The architecture of UniMamba is shown in Fig. 2. Following [48], we use an encoder-decoder architecture to stack UniMamba Blocks for further hierarchical feature extraction. As a general voxel-based 3D backbone, we replace the backbone in LION [24] to construct our detector. For Argoverse 2 [41], we build on SAFDNet [49]. The configuration of the detection head and loss functions remains consistent with the baseline.
@@ 4.1 Datasets and Metrics
nuScenes [2] is a sophisticated outdoor dataset comprising 1,000 scenarios, with 700 allocated for training, 150 for validation, and 150 for testing. It encompasses a diverse array of object annotations across 10 categories, representing the majority of traffic participants. The size of objects varies significantly among these categories, which poses considerable challenges for detection. We evaluate the effectiveness of our approach using the provided assessment tools, and reporting metrics including NDS and mAP.
@@ 4.2 Implementation Details
We implement our UniMamba based on LION [24]. Following previous methods, we use voxel sizes of $(0.3m,0.3m,0.25m)$, $(0.32m,0.32m,0.1875m)$, and $(0.4m,0.4m,0.25m)$ for voxelization of the nuScenes, Waymo, and Argoverse 2 datasets, respectively, where the number of voxel feature channels is set to 128. The downsampling stride for encoder-decoder is set to $\{1,2,2\}$, with each scale containing 1 layer of UniMamba Block. For each UniMamba block, the number of channel groups $M$ is set to 4, with $J=2$ groups applying GSE and the remaining two applying LSE. The other training hyperparameters are configured consistently with [24]. All experiments are conducted using 8 Tesla A800 GPUs, utilizing the AdamW [26] optimizer. For more network details, please refer to the appendix.
@@ 4.4 Ablation Study
Effect of Spatial Locality Modeling. To validate the effectiveness of the proposed SLM in mitigating the loss of spatial locality, we compare it with several popular space-filling curves. The ”Random” refers to the configuration where no spatial curve is applied. As shown in rows 2-4 of Tab. 4, when neither SLM nor any spatial curve is used, performance drops significantly (- 1.4 mAP), demonstrating the importance of spatial locality. Compared with the standard Z-order curve, the Hilbert curve which better preserves locality shows a 0.2 mAP advantage. With SLM, the ”Random” achieves 67.4 mAP, comparable to the performance with space-filling curves. Additionally, the Z-order curve achieves performance parity with the Hilbert curve (67.6 mAP) but requires only 1/17 time cost. These results demonstrate the effectiveness of Spatial Locality Modeling in capturing the dynamic structure embedding.